OpenStudy (anonymous):

find the first four partial sums k=1sigma tends to infinity (2/(5^k)-1)

OpenStudy (anonymous):

is this \[\frac{2}{5^k-1}\]

OpenStudy (anonymous):

yes yes..

OpenStudy (anonymous):

replace \(k\) by 1 get \[\frac{2}{5^1-1}=\frac{2}{4}=\frac{1}{2}\]so that is your first partial sum

OpenStudy (anonymous):

replace \(k\) by 2 get \[\frac{2}{5^2-1}=\frac{1}{12}\] so your second partial is \[\frac{1}{2}+\frac{1}{12}=\frac{7}{12}\]

OpenStudy (anonymous):

sory sory 5^(k-1)

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

replace \(k\) by 1 get \[\frac{2}{5^0}=2\] so that is the first partial

OpenStudy (anonymous):

now replace \(k\) by 2 and get \[\frac{2}{5^1}=\frac{2}{5}\] second partial is \[2+\frac{2}{5}=\frac{12}{5}\]

OpenStudy (anonymous):

etc etc

OpenStudy (anonymous):

i.e. replace k by successive integers, and add

OpenStudy (anonymous):

thnks. how can i find the series closed form of nth partial sums and how can i find this series in converge or diverge

OpenStudy (anonymous):

plz help me.;(